OpenStudy (anonymous):

The sum of Emil’s and Rosa’s ages is 27 years. Nine years ago Emil was twice as old as Rosa was then. How old is each person? What are the two equation ?

OpenStudy (anonymous):

x+y=27 is the first equation right ?

jigglypuff314 (jigglypuff314):

but then it says "nine years ago" so 27 - 9 - 9 = then x + y = (27- 9 -9)

jigglypuff314 (jigglypuff314):

then Emil was twice as old as Rosa so x = 2y plug into x + y = 9 2y + y = 9 so y = 3 and x = 6 then add back 9 to both to get how old they are now

OpenStudy (anonymous):

so is x+y=27 the right question

OpenStudy (anonymous):

Sorry I still don't get it

OpenStudy (anonymous):

So I just want two equation that it

jigglypuff314 (jigglypuff314):

you can get two equations from the "nine years ago" equations are x+y = 9 and x = 2y then once you get the x and y values present time : (x+9) and (y+9)

OpenStudy (anonymous):

u sure it x=2y

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (skullpatrol):

x-9 = 2(y-9)

OpenStudy (skullpatrol):

Nine years ago Emil was twice as old as Rosa was then.